OpenStudy (anonymous):

what is the sum of y2 and 5y2

OpenStudy (anonymous):

6y2

OpenStudy (anonymous):

Pre algebra join the room physical education :)

OpenStudy (anonymous):

if its tough to understand, u can always "factor"\[y^2+5y^2=y^2(1+5)=6y^2\]